OpenStudy (amonoconnor):

Is this the correct answer? "The formula for the surface area or a sphere is: SA = 4pi*r^2. If a snowball melts so that its SA decreases at a rate of 1cm^2/min, find the rate at which the radius decreases when the diameter is 1 meter. Any and all help is greatly appreciated!

OpenStudy (misty1212):

HI!! ( again )

OpenStudy (misty1212):

\[S=4\pi r^2\\ S'=8\pi rr'\]

OpenStudy (amonoconnor):

\[\frac{dr}{dt} = \frac{1}{400\pi} cm/\min\]

OpenStudy (misty1212):

your are told \(S'=1\) put \[1=8\pi rr'\] solve for \(r'\)

OpenStudy (misty1212):

if the diameter is \(1\) then the radius is \(\frac{1}{2}\)

OpenStudy (misty1212):

so no, i don't think that answer is correct

OpenStudy (misty1212):

oooh i see the units have changed doe

OpenStudy (misty1212):

very tricky

OpenStudy (misty1212):

ok now your answer looks a lot better

OpenStudy (misty1212):

\[1=8\times 50\pi r'\] so yeah

OpenStudy (amonoconnor):

Here's what I did: \[SA = 4\pi*r^2\] \[\frac{dSA}{dt} = 4\pi*2r*\frac{dr}{dt}\] \[(1cm^2/\min) = 4\pi(2*50cm)*\frac{dr}{}\] \[1 = 400\pi*\frac{dr}{dt}\] \[\frac{dr}{dt} = \frac{ 1 }{400\pi}\]

OpenStudy (amonoconnor):

Is that wrong? :/

OpenStudy (misty1212):

yeah right lot easier to write \(r'\) instead of \(\frac{dr}{dt}\) but yes you are correct i did not notice they changed the units from cm to mm very tricky (well not really, if i knew how to read carefully) you are right as far as i can see

OpenStudy (amonoconnor):

Thanks again misty;)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]